OpenStudy (anonymous):

A cylinder with radius and height 2r +4 contains a cube with edge length r When it's a good idea to start r√2. What fraction of the cylinders volume is taken up by the cube. Use a rational expression and write your answer in simplified form.

OpenStudy (anonymous):

what does the second sentence mean? "When it's a good idea to start \[r*\sqrt{2}\]

OpenStudy (anonymous):

Sorry I messed the question up when I was trying to put in a square symbol. "When's it a good idea to start" shouldn't be in there.

OpenStudy (anonymous):

? So for this question can't you just calculate both volumes set up the ratio?

OpenStudy (anonymous):

I am not sure how to

OpenStudy (anonymous):

ok lets go through it. Volume of a cube is side^3, and this cube has side of length root(2)*r

OpenStudy (anonymous):

\[\left( \sqrt{2}*r \right)^3\]

OpenStudy (anonymous):

then the cylinder has a volume of length*base.

OpenStudy (anonymous):

height*

OpenStudy (anonymous):

so what is the area of the base?

OpenStudy (anonymous):

it would be the area of a circle with radius 2r+4

OpenStudy (anonymous):

can you calculate it?

OpenStudy (anonymous):

A = 2r +4 * 2r

OpenStudy (anonymous):

Oh I know whats confusing you. The problem writers were kind of mean when they used r as the variable, maybe it'll help if you consider the "r" an "x" in this case, because its not REALLY the radius.

OpenStudy (anonymous):

so the radius of the base is 2x+4 thus the area is 2pi(radius)^2

OpenStudy (anonymous):

so you're looking at \[2\pi*\left( 2x+4 \right)^2\]

OpenStudy (anonymous):

^^ that is the area of the base

OpenStudy (anonymous):

so if you have the base, then you multiply by the height to get volume.

OpenStudy (anonymous):

are you still with me?

OpenStudy (anonymous):

:( I hope you got it